A valtage of 200v is connected across a capacitor of a capacitance of 2.5uF. calculate the charge deposited on each plate


1
Expert's answer
2021-07-30T03:54:02-0400

Energy stored in "2.5\ \mu f" capacitor ;


"U_{2.5}=0.5\times C\times V^2\\U_{2. 5}=0.5×2.5×10 ^{-6}×200^2"


"U_{2.5}=0.05J"